# Non-canonical 3-form dark energy

### v1

@@ I Driving the late-time acceleration with a 3-form
One such alternative relies on 3-form fields Koivisto and Nunes (2009, 2013). These have been shown to yield accelerated solutions in cosmology, useful to explain both early (primordial inflation) and late (dark energy) time epochs of our universe Koivisto and Nunes (2010). Inflation under 3-form fields has been further studied in Ref. Germani and Kehagias (2009); Mulryne et al. (2012); De Felice et al. (2012); Kumar et al. (2014); Barros and Nunes (2016); Sravan Kumar et al. (2016). More generally, they are also useful to describe the geometry of stars, black holes, wormholes or singularities Barros and Lobo (2018); Barros et al. (2020); Bouhmadi-López et al. (2021a); Barros et al. (2021); Bouhmadi-López et al. (2021b); Morais et al. (2017); Bouhmadi-López et al. (2018), as well as Bianchi cosmologies Normann et al. (2018). The general formalism of 3-forms has been developed in previous studies such as those in Refs. Koivisto and Nunes (2010); Wongjun (2017); Beltrán Almeida et al. (2020).
@@ II The Lagrangian
where the factor $2$ simplifies the equations from now on.
In the following two sections, we will particularize this Lagrangian to perform the analysis of the dynamical system (derived from the main cosmological equations) and the corresponding 3-form cosmology. Our main goal is to consider a 3-form dark energy Lagrangian with a dust-like energy density, ${\gamma=1}$, which eventually freezes. The acceleration of the universe at late times is achieved by choosing an exponential form for the interaction term.
@@ III Dynamical system analysis
To investigate the stability of the system, we study the trajectories in the vicinity of the fixed points. Small perturbations are considered around them in order to define the stability matrix, from which the eigenvalues, ${(e_{1},e_{2},e_{3})}$ listed in Appendix B, are derived. These values allow for the determination of the stability of the critical points, which can then be classified into three distinct categories: stable, unstable, or saddle points. We find that the behavior of the dynamical system is contingent upon the values of the 3-form potential slope.
@@ V Conclusion
In this work, we explored the suggestions of Ref Wongjun (2017), introducing a generalized solution for describing a 3-form field A that behaves as a cosmological fluid with a constant equation of state parameter. We also developed a viable cosmological model by formulating a Lagrangian that endows the 3-form with the dynamical properties of both dust at the beginning, and dark energy accelerating the universe in late times. The acceleration is achieved by including a self-interacting exponential term of the form ${V\propto\exp(-3A^{2}/\lambda)}$, which stabilizes the solutions in an inflationary state. Resorting to dynamical systems techniques, we thoroughly analyzed the solutions considering a flat FLRW background universe filled with matter and the 3-form fluid. The latter, due to homogeneity and isotropy, can be fully parameterised in terms of a scalar, $\chi$, function of cosmic time only.
@@
For future work, it would be worthwhile to test the model against observational data. In particular, datasets of distance measurements, e.g., of supernovae and baryonic acoustic oscillations, could have the potential to constrain the perimeter of the model, which is subject to the usual fine-tuning of dark energy models, without having to go into the non-linear regime.

### v2

@@ I Driving the late-time acceleration with a 3-form
One such alternative relies on 3-form fields Koivisto and Nunes (2009, 2013). These have been shown to yield accelerated solutions in cosmology, useful to explain both early (primordial inflation) and late (dark energy) time epochs of our universe Koivisto and Nunes (2010); Morais et al. (2017a); Bouhmadi-López et al. (2017). Inflation under 3-form fields has been further studied in Ref. Bouhmadi-López et al. (2017); Germani and Kehagias (2009); Mulryne et al. (2012); De Felice et al. (2012); Kumar et al. (2014); Barros and Nunes (2016); Sravan Kumar et al. (2016). More generally, they are also useful to describe the geometry of stars, black holes, wormholes or singularities Barros and Lobo (2018); Barros et al. (2020); Bouhmadi-López et al. (2021a); Barros et al. (2021); Bouhmadi-López et al. (2021b); Morais et al. (2017b); Bouhmadi-López et al. (2018), as well as Bianchi cosmologies Normann et al. (2018) or extra-dimensional branes Barros and Beltrán Jiménez (2024). The general formalism of 3-forms has been developed in previous studies such as those in Refs. Koivisto and Nunes (2010); Wongjun (2017); Beltrán Almeida et al. (2020).
@@ II The Lagrangian
where the factor $2$ simplifies the equations from now on. The form of the Lagrangian (11) derives from a phenomenological approach, but the action can be framed in an effective field theory. The latter would be valid in some low-energy range of a more fundamental high-energy model, where non-canonical 3-form fluxes can emerge after dimensionality reduction Cremmer et al. (1978); Avetisyan et al. (2022); Evnin et al. (2024); Vanichchapongjaroen (2021); Nitta and Yokokura (2018).
In the following two sections, we will particularize this Lagrangian to perform the analysis of the dynamical system (derived from the main cosmological equations) and the corresponding 3-form cosmology. Our main goal is to consider a 3-form dark energy Lagrangian with a dust-like energy density, ${\gamma=1}$, which eventually freezes. The acceleration of the universe at late times is achieved by choosing an exponential form for the interaction term. Although this choice is again phenomenological, we consider self-interactions of more fundamental fields. In superstring cosmology, it has also been shown that an exponential potential can stabilize the dilaton Barreiro et al. (1998).
@@ III Dynamical system analysis
To investigate the stability of the system, we study the trajectories in the vicinity of the fixed points. Small perturbations are considered around them in order to define the stability matrix, from which the eigenvalues, ${(e_{1},e_{2},e_{3})}$ listed in Appendix A, are derived. These values allow for the determination of the stability of the critical points, which can then be classified into three distinct categories: stable, unstable, or saddle points. We find that the behavior of the dynamical system is contingent upon the values of the 3-form potential slope.
@@ V Conclusion
In this work, we explored the suggestions of Ref Wongjun (2017), introducing a generalized solution for describing a 3-form field A that behaves as a cosmological fluid with a constant equation of state parameter. We also developed a viable cosmological model by formulating a Lagrangian that endows the 3-form with the dynamical properties of both dust at the beginning, and dark energy accelerating the universe in late times. The acceleration is achieved by including a self-interacting exponential term of the form ${V\propto\exp(-3A^{2}/\lambda)}$, which stabilizes the solutions in an inflationary state. Resorting to dynamical systems techniques, we thoroughly analyzed the solutions considering a flat FLRW background universe filled with matter and the 3-form fluid. The latter, due to homogeneity and isotropy, can be fully parametrized in terms of a scalar, $\chi$, function of cosmic time only.
@@
For future work, it would be worthwhile to test the model against observational data. In particular, data sets of distance measurements, e.g. of supernovae and baryonic acoustic oscillations, could have the potential to constrain the model, which is subject to the usual fine-tuning of dark energy models, without having to go to the perturbation level. However, the latter will be necessary to test the model with cosmic microwave background and large-scale structure observables, probing high and low redshifts, respectively.